slove dy/dx= x-e^x
1
Expert's answer
2011-12-15T07:11:35-0500
to solve this we rewrite it in the form dy=(x-e^x)dx
then we can integrate
both sides and we get
integral 1*dy=integral (x-e^x)dx
so y=0.5*x^2-e^x+C,
where C - constant is real number